OpenStudy (anonymous):

If a = -12/5, then 5x + 2y = 6 and 3x - ay = 4 are parallel. How do I solve this?

OpenStudy (destiny23):

What can you say about these two lines having equations: 2x + 3y = 9 And, 4x + 6y = 18

OpenStudy (mayankdevnani):

any guesses ??

OpenStudy (mayankdevnani):

@brimae97

OpenStudy (anonymous):

Im not sure..

Parth (parthkohli):

Solve what?

Parth (parthkohli):

Is this a true/false problem?

OpenStudy (anonymous):

It's a true or False question but i just want to know how i can determine how to figure it out.

OpenStudy (mayankdevnani):

Remember :- Consider two lines :- \[\large \bf a_1x+b_1y+c_1=0~and~a_2x+b_2y+c_2=0\]

OpenStudy (mayankdevnani):

If they are parallel,then use this :- \[\large \bf \frac{a_1}{a_2}=\frac{b_1}{b_2}=\frac{c_1}{c_2}\]

OpenStudy (mayankdevnani):

^^^^^ `This is the condition for parallel lines`

Parth (parthkohli):

Hmm, the second \(=\) should be \(\ne\)

OpenStudy (anonymous):

The answer is that the equations aren't parallel so it's false?

OpenStudy (mayankdevnani):

with the help of above equation ,find out that they are parallel or not ??

OpenStudy (mayankdevnani):

Show your work @brimae97

Parth (parthkohli):

That condition you said is for coincident lines. For parallel lines, you need to have no solutions.

OpenStudy (mayankdevnani):

yeah !! AGAIN SORRY !

OpenStudy (mayankdevnani):

lolz

OpenStudy (mayankdevnani):

`My DULL mind`

OpenStudy (mayankdevnani):

I think my mind is on HOLIDAY

OpenStudy (mayankdevnani):

xD